Suppose the total demand function for a good was made up of 10 identical, individual demanders all with a demand function given by: 10 0.2P Find the inverse, total demand function and fill in the missing values below. P 50 (Round to the nearest two decimal places if necessary.)
Answer #1

Demand function is given by qD = 10 - 0.2P. Market demand is 10qD = QD = 100 - 2P.

Inverse demand is 2P = 100 - QD or P = 100/2 - QD/2

Hence inverse demand is P = 50 - 0.5QD

Answer to the box is 0.5